# Frustration-free free fermions and beyond

### v1

@@ IV.1 Derivation of BdG form
$\hat{H}_{\bm{R}}\coloneqq\hat{\bm{\gamma}}_{\bm{R}}^{\top}A_{\bm{R}}\hat{\bm{\gamma}}_{\bm{R}}+C_{\bm{R}},$ (74)
@@
$\hat{c}_{\bm{R}\tau}\coloneqq\frac{1}{2}\left(\hat{\gamma}_{\bm{R}\tau 1}+{\mathrm{i}}\hat{\gamma}_{\bm{R}\tau 2}\right).$ (75)
@@
$\hat{\Psi}_{\bm{R}}=\frac{1}{2}\begin{pmatrix}\mathbbm{1}&{\mathrm{i}}\mathbbm{1}\\ \mathbbm{1}&-{\mathrm{i}}\mathbbm{1}\end{pmatrix}\hat{\bm{\gamma}}_{\bm{R}},$ (76)
$h_{\bm{R}}=\begin{pmatrix}\mathbbm{1}&{\mathrm{i}}\mathbbm{1}\\ \mathbbm{1}&-{\mathrm{i}}\mathbbm{1}\end{pmatrix}{\mathrm{i}}A_{\bm{R}}\begin{pmatrix}\mathbbm{1}&\mathbbm{1}\\ -{\mathrm{i}}\mathbbm{1}&{\mathrm{i}}\mathbbm{1}\end{pmatrix}.$ (77)
@@
$=\begin{pmatrix}O&\mathbbm{1}\\ \mathbbm{1}&O\end{pmatrix}\begin{pmatrix}\mathbbm{1}&-{\mathrm{i}}\mathbbm{1}\\ \mathbbm{1}&{\mathrm{i}}\mathbbm{1}\end{pmatrix}{\mathrm{i}}A_{\bm{R}}^{\top}\begin{pmatrix}\mathbbm{1}&\mathbbm{1}\\ {\mathrm{i}}\mathbbm{1}&-{\mathrm{i}}\mathbbm{1}\end{pmatrix}\begin{pmatrix}O&\mathbbm{1}\\ \mathbbm{1}&O\end{pmatrix}$
$=\begin{pmatrix}\mathbbm{1}&{\mathrm{i}}\mathbbm{1}\\ \mathbbm{1}&-{\mathrm{i}}\mathbbm{1}\end{pmatrix}(-{\mathrm{i}}A_{\bm{R}})\begin{pmatrix}\mathbbm{1}&\mathbbm{1}\\ -{\mathrm{i}}\mathbbm{1}&{\mathrm{i}}\mathbbm{1}\end{pmatrix}=-h_{\bm{R}}.$ (78)
@@ IV.2 Frustration-free condition
Now, we are ready to derive the frustration-free condition. Let $\bm{\psi}_{\bm{R}\alpha}$ $(\alpha=1,2,\cdots,A_{\bm{R}})$ be eigenvectors of $h_{\bm{R}}$ with positive eigenvalues $\mu_{\bm{R}\alpha}/2>0$. Due to the symmetry relation (73), $\bm{\phi}_{\bm{R}\alpha}\coloneqq U_{\mathcal{P}}\bm{\psi}^{*}_{\bm{R}\alpha}$ is an eigenvector with a negative eigenvalue $-\mu_{\bm{R}\alpha}/2<0$; that is,
@@
Consider a state given by |$\ket{\Phi_{\bm{R}}}\coloneqq\prod_{\alpha=1}^{A_{\bm{R}}}\hat{\psi}_{\bm{R}\alpha}\ket{0}$, which satisfies $\hat{\psi}_{\bm{R}\alpha}\ket{\Phi_{\bm{R}}}=0$ for all $\alpha$. Hence, |$\ket{\Phi_{\bm{R}}}$ is a ground state of $\hat{H}_{\bm{R}}$ and satisfies $\hat{H}_{\bm{R}}\ket{\Phi_{\bm{R}}}=0$. Following the same discussion as the U(1)-symmetric case, we arrive at the frustration-free condition
@@ IV.3.1 The Kitaev chain
$\hat{H}_{i}\coloneqq\frac{t_{0}}{2}+\frac{1}{2}\hat{\Psi}_{i}^{\top}\begin{pmatrix}\frac{\mu}{2}&-t&0&-\Delta\\ -t&\frac{\mu}{2}&\Delta&0\\ 0&\Delta&-\frac{\mu}{2}&t\\ -\Delta&0&t&-\frac{\mu}{2}\\ \end{pmatrix}\hat{\Psi}_{i}$ (94)
@@
$\hat{\psi}_{i}=\frac{\cos\theta}{\sqrt{2}}(\hat{c}_{i}-\hat{c}_{i+1})+\frac{\sin\theta}{\sqrt{2}}(\hat{c}_{i}^{\dagger}+\hat{c}_{i+1}^{\dagger}).$ (95)

### v2

@@ IV.1 Derivation of BdG form
$\hat{H}_{\bm{R}}\coloneqq i\hat{\bm{\gamma}}_{\bm{R}}^{\top}A_{\bm{R}}\hat{\bm{\gamma}}_{\bm{R}}+C_{\bm{R}},$ (77)
@@
$\hat{c}_{\bm{R}\tau}\coloneqq\frac{1}{2}\left(\hat{\gamma}_{\bm{R}\tau 1}+i\hat{\gamma}_{\bm{R}\tau 2}\right).$ (78)
@@
$\hat{\Psi}_{\bm{R}}=\frac{1}{2}\begin{pmatrix}\mathbbm{1}&i\mathbbm{1}\\ \mathbbm{1}&-i\mathbbm{1}\end{pmatrix}\hat{\bm{\gamma}}_{\bm{R}},$ (79)
$h_{\bm{R}}=\begin{pmatrix}\mathbbm{1}&i\mathbbm{1}\\ \mathbbm{1}&-i\mathbbm{1}\end{pmatrix}iA_{\bm{R}}\begin{pmatrix}\mathbbm{1}&\mathbbm{1}\\ -i\mathbbm{1}&i\mathbbm{1}\end{pmatrix}$. (80)
@@
$=\begin{pmatrix}O&\mathbbm{1}\\ \mathbbm{1}&O\end{pmatrix}\begin{pmatrix}\mathbbm{1}&-i\mathbbm{1}\\ \mathbbm{1}&i\mathbbm{1}\end{pmatrix}iA_{\bm{R}}^{\top}\begin{pmatrix}\mathbbm{1}&\mathbbm{1}\\ i\mathbbm{1}&-i\mathbbm{1}\end{pmatrix}\begin{pmatrix}O&\mathbbm{1}\\ \mathbbm{1}&O\end{pmatrix}$
$=\begin{pmatrix}\mathbbm{1}&i\mathbbm{1}\\ \mathbbm{1}&-i\mathbbm{1}\end{pmatrix}(-iA_{\bm{R}})\begin{pmatrix}\mathbbm{1}&\mathbbm{1}\\ -i\mathbbm{1}&i\mathbbm{1}\end{pmatrix}=-h_{\bm{R}}.$ (81)
@@ IV.2 Frustration-free condition
Now, we are ready to derive the frustration-free condition. Let $\bm{\psi}_{\bm{R}\alpha}$ $(\alpha=1,2,\cdots,A_{\bm{R}})$ be eigenvectors of $h_{\bm{R}}$ with positive eigenvalues $\mu_{\bm{R}\alpha}/2>0$. Due to the symmetry relation (76), $\bm{\phi}_{\bm{R}\alpha}\coloneqq U_{\mathcal{P}}\bm{\psi}^{*}_{\bm{R}\alpha}$ is an eigenvector with a negative eigenvalue $\mu_{\bm{R}\alpha}/2<0$; that is,
@@
[…]
@@ IV.3.1 The Kitaev chain
$\hat{H}_{i}\coloneqq\frac{t_{0}}{2}+\frac{1}{2}\hat{\Psi}_{i}^{\dagger}\begin{pmatrix}\frac{\mu}{2}&-t&0&-\Delta\\ -t&\frac{\mu}{2}&\Delta&0\\ 0&\Delta&-\frac{\mu}{2}&t\\ -\Delta&0&t&-\frac{\mu}{2}\\ \end{pmatrix}\hat{\Psi}_{i}$ (97)
@@
$\hat{\psi}_{i}=\frac{\cos\theta}{\sqrt{2}}(\hat{c}_{i}-\hat{c}_{i+1})+\frac{\sin\theta}{\sqrt{2}}(\hat{c}_{i}^{\dagger}+\hat{c}_{i+1}^{\dagger})$ (98)